OpenStudy (anonymous):

Tammy is 14 years old she has a younger sister and an older brother her sister is 12 years old the mean of all their ages is 14 how old is Tammy brother ?

OpenStudy (anonymous):

Lets change the numbers so you can work another example. Let's say Tammy is 16 years old and her younger sister is 13. Let's say the mean of all three ages is 16. How old is her brother? Well, from the definition of the mean, we know that (13 + 16 + x) / 3 = 16, where x is her brother's age. Now use arithmetic and solve for x. Multiply each side by 3 to get 13 + 16 + x = 48. Subtract 13 and 16 from each side to get x = 19. So for this example her brother's age is 19. Now change the numbers to the ones you provided.

OpenStudy (anonymous):

so how old is the brother

OpenStudy (anonymous):

All right well lets wlk through this together. Do you know how to find the variable?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay. We will do the example so you can figure the real answer on your own. First we have to solve the parenthesis. So what is 13 + 16?

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

sorry back

OpenStudy (anonymous):

thats ok.

OpenStudy (anonymous):

29

OpenStudy (anonymous):

Good. Now the equation will look like this. 29 + x / 3 = 16. The next step is to divide 29 and 3. This is going to become a decimal so just round to the nearest whole number.

OpenStudy (anonymous):

9

OpenStudy (the_fizicx99):

It's really not that hard, simply work backwards... To find the mean you add all the number and divide by how many there are, so there's 3. 14 + 12 + x / 3 = 14 You want to find x which is their brother so you cancel out the division by multiplying both sides by 3.  Multiply both sides by 3, So you have 14 + 12 + x = 42 14 + 12 = 26 26 + x = 42 -26 -26 x = 16 So their brother's age is 16. Check my plugging 16 for x and dividing by 3 14 + 12 + 16/3 42/3 = 14 <-- 14 is the original mean.